OpenStudy (anonymous):

Find the difference. 1/x^2+5x (minus) 4/x^2-25

OpenStudy (callisto):

\[ \frac{1}{x^2+5x}-\frac{ 4}{x^2-25}\]\[=\frac{1}{x(x+5)}-\frac{ 4}{(x-5)(x+5)}\]\[=\frac{(x-5)}{x(x+5)(x-5)}-\frac{ 4x}{x(x-5)(x+5)}\]\[=\frac{x-5- 4x}{x(x-5)(x+5)}\]\[=\frac{-5- 3x}{x(x-5)(x+5)}\]\[=\frac{5+3x}{x(5-x)(x+5)}\]

OpenStudy (anonymous):

Is this simplest form?

OpenStudy (callisto):

What do you think?

OpenStudy (anonymous):

I don't know! lol

OpenStudy (callisto):

It may be :)

OpenStudy (anonymous):

lol okay, well I have more questions. So please help. lol